solve the system by the substitution method
6x-3y=3
10x=y+7

Answer by mananth(16949)

6x-3y=3
10x=y+7
x= y+7/10
substitute the value of x in 6x-3y=3
6((y+7)/10-3y=3
3/5(y+7)-3y=3
multiply by 5
3(y+7)-15y=15
3y+21-15y=15
-12y=-6
divide by -12
y=1/2
plug value of y in the equation 10x=y+7
10x=1/2+7
10x=15/2
divide by 10
x=15/2 *1/10
x= 3/4